Rick always does an awesome job I think he deserves a raise
Replaced the electrical disconnect and the temp control on the walkin freezer. Verified operation after repairs. Set the defrost clock to the correct time after having the unit off for repairs. Repairs were approved by Harry Frazier.
Checked the walkin freezer for not maintaining temp after walking Harry thru bypassing the failed disconnect on Saturday night. Found the temp 20 deg with the condenser cycled off. Found the temp control off by 15 deg. Set the temp to 15 below to maintain a box temp of 0 deg. Will return to replace the disconnect and the temp control.
Checked the walkin cooler for too warm complaint. Found the condenser off on arrival. Started the unit and found the system pulling down to 40 deg. Added 7.5 oz of R-22 to the 2# 4oz required charge. System pulled down to 35 deg after adding refrigerant.
Replaced the headmaster head pressure control on the walkin cooler condenser for proper head pressure control during low ambient conditions. Verified operation after repairs. Repairs were approved by Harry Frazier.
Checked the walkin cooler for the condenser not running. Found the unit off at the disconnect on arrival. Restored the power and checked the operation. Found the liquid pressure excessively high due to a failed headmaster control. Disabled the headmaster valve. Verified the operation and liquid pressure after disabling the valve. Will return to replace the valve before cold weather. System is operating normal with a box temp of 35 deg.
Checked the walk in cooler for no cool after replacing the expansion valve power head from the previous day. Found the expansion valve failed shut with the condenser pumped down and cycled off. Removed the refrigerant, replaced the expansion valve, liquid dryer and low pressure switch. Evacuated the system, added the removed refrigerant, started the unit and verified operation. Ajusted the expansion valve to the proper superheat. Found the unit cooling normal with the box temp dropping to 37 deg from the 50 deg start up temp.
9/25/19--Checked the walkin cooler for too warm. Found the unit pumped down due to a failed expansion valve power head. Made a temporary repair by placing a shim under the head to open the valve so the cooler would not have to be unloaded. Called the shop and had a replacement head picked up. Will return tomorrow to replace the head. 9/26/19--Replaced the power head. Verified operation after repairs. Repairs were approved by Harry Frazier.
Checked the mitsubishi mini split for a P5 condensate backup alarm on the thermostat. Cleaned the condensate pump sump to remove the sludge blockage preventing condensate removal. Verified operation after cleaning and resetting the alarm. Showed the maintenance man on site how to clean and maintain the condensate assembly.
Checked the 3.5 ton mini split for the kitchen for no cool. Found the unit low on charge. Leak searched the system. Found the leak at the roof deck penetration above the pantry hard ceiling. Repaired the insulation that was removed for the leak search. Removed and weigthed the existing refrigerant for the correct weight in recharge. Started the system and verified operation. Zach will contact the customer with the details on replacing the line set due to the location of the leak. Warranty.
Checked the walkin cooler and freezer for not maintaining temp. Found the cooler low on charge. Added R-22 to the proper charge. Found the freezer also low on charge. Added R-404a to the proper charge. Verified operation of both after repairs. Repairs were approved by Harry Frazer.
Checked the Mitsubishi split for the kitchen for no cool. Found the system excessively low on charge due to a refrigerant leak at the suction line flare connection at the condenser. Reflared the connection, evacuated the system, added r410a, leak checked the repairs, started the unit and verified operation. Ran out of R410a and shorted the charge by 1 lb 10 oz. Will return to finish the charge tomorrow morning. Warranty. 8/29/19--Added 1 lb 10 oz. to the correct charge. Verified operation. Warranty.
6/6/19--Checked the 15 ton Goodman Rtu for the dining room for not keeping up. Found two of the three condenser fans not running with the compressors cycling on high pressure. Found the center motor with open windings, the East motor with seized bearings and the West motor running with worn bearings. Drilled holes at the bearing housings on the East motor and freed the bearings with penetrating oil for temporary repairs. Was able to get two of the three motors running till the motors can be replaced tomorrow. Blocked the center fan housing to stop the condensing air from bypassing thru the failed motor. Unit is cooling normal with two of the three motors running. Also checked the 18,000 btu goodman split for Liz's West office for no cool. Found the system low on charge. Added 2# R-22 to the correct charge. Will return tomorrow to replace the three condenser fan motors when the motors can be picked up and delivered. 6/7/19--Replaced the three condenser fan motors and capacitors in the G
Checked the ten year old LG 3.5 ton mini split for the kitchen for not maintaining temp after adding 2# of refrigerant on 4/22/19. Found the system low on charge. Leak searched the system. Found the leak in the indoor cassette evaporator. Recommended replacing the system do to the age and condition. Sent Zach the info for quoting.
Added low ambient controls to the two ton split for the Kitchen too prevent evaporator freeze ups. Verified operation after repairs. Repairs were approved by Harry Frazier.
Reassembled the ceiling cassette for the 3.5 ton LG mini-split after the maintenance man cleaned the plugged evaporator. Started the system and checked the operation. Added R410a to the proper charge. Evaporator split was 20 deg after repairs. Repairs were approved by Harry Frazier.